At the homecoming football game, the senior class officers sold slices of pizza for $0.75 each and hamburgers for $1.35 each. They sold 40 more slices of pizza than hamburgers, and sales totaled $292.50. How many slices of pizza did they sell?
Answer by checkley77(12569)

p=h+40
.75p+1.35h=292.50
.75(h+40)+1.35h=292.50
.75h+30+1.35h=292.50
2.1h=292.50-30
2.1h=262.5
h=262.5/2.1125
h=125 hamburgers were sold.
p=125+40=165 pieces of pizza were sold.
Proof:
.75*165+1.35*125=292.50
123.75+168.75=292.50
292.50=292.50